# An Overview and Solution for Democratizing AI Workflows at the Network Edge

### v2

@@ 7 Evaluation Methodology
To evaluate the proposed NAOMI solution we compare it to the current release of the O-RAN AI/ML Framework each consisting of the components listed in Table 8. The first column lists different AI/ML components and the other two columns list the open-source tools utilized by the compared solutions: NAOMI and O-RAN. The listed tools were presented and analyzed in Section 5.
In this section, we first elaborate on the experimental infrastructure and setup, summarized in Table 9, then we present the methodology used for evaluating NAOMI and comparing both solutions. Further, we identify any limitations to the evaluation and results.
@@ 7.2 Deployment Time
NAOMI supports deployment on a distributed setup, as described in Section 7.1. To test the deployment time of NAOMI on a distributed setup, we automated the deployment using an Ansible playbook. We tested the deployment time from the time the Ansible playbook started executing to the time all the nodes and NAOMI deployments were in a ready state, providing insight into the overhead of a distributed system.
@@ 7.3 Resource Usage
During the example AI/ML workflow execution, resource utilization was measured at 1 s intervals and CPU usage and RAM usage graphs for both solutions were plotted with timestamps indicating the start of each of the workflow stages.
@@ 7.5 Distributed Training
Batch size affects the performance of the distributed training [111], so training time was measured for different batch sizes and compared between different scenarios listed in Section 7.1. For a comparison, we included training time on O-RAN AI/ML Framework with the default scenario. Resource usage metrics and training time were measured in the same manner as described in Section 7.3.
@@ 7.6 Inference Latency
As NAOMI supports deployment on a distributed setup, described in Section 7.1, the inference was evaluated for NAOMI when deployed on such setup. We evaluate the ability of load-balancing on multiple nodes and any improvements in mean latency as inference is performed distributively across more CPU cores. On a distributed setup there are 16 cores available for model inference compared to 4 cores on a centralized setup. With this modification, a theoretical speedup of up to 4 times (mean latencies decrease by a factor of 4) is expected to be achieved.
@@ 8.2 QoE AI/ML Workflow Execution
Further, the O-RAN solution did not complete the execution of the workflow with the $100x$ dataset, presumably due to timeouts in data extraction stages, which were observed in the logs. NAOMI completed the execution in $1095s$ with $100x$ dataset size, which is approximately 3 times longer than the $10x$ dataset. The difference in execution times for NAOMI between the $1x$ and $10x$ datasets is 75 seconds, and between the $10x$ and $100x$ datasets, it is 736 seconds. This indicates that the portion of the workflow execution that takes longer as the dataset size grows increases roughly linearly with the size of the dataset.
@@ 8.3 Distributed Training
From Fig. 11, we can see that NAOMI on a centralized setup executed the QoE workflow in $284s$, $359s$, and $1095s$ for $1x$, $10x$, and $100x$ dataset respectively, while NAOMI on a distributed setup executed the QoE workflow in $356s$, $825s$, and $5435s$ for described dataset sizes. For model training the distributed setup together with distributed model training performs significantly worse. The centralized setup performed $31\%$ faster for $1x$ dataset size, $68\%$ for $10x$ and $83\%$ for $100x$. This shows a significant overhead of distributed model training for this set of training parameters. The distributed setup does not affect data extraction, model deployment, and other (overhead) tasks. In all cases, both setups performed these tasks with a difference of up to 3 seconds, which is within the margin of error.
@@ 8.5 Workflows Beyond O-RAN
QoE workflow example with 284 seconds takes, on average, the least time to complete. ImClass and VWW take 670s and 732s respectively, while the MNIST workflow example takes 1340 seconds.

### v3

@@ 7 Evaluation Methodology
To evaluate the proposed NAOMI solution we compare it to the current release of the O-RAN AI/ML Framework each consisting of the components listed in Table 8. The first column lists different AI/ML components and the other two columns list the open-source tools utilized by the compared solutions: NAOMI and O-RAN. The listed tools were presented and analyzed in Section 5. In this section, we first elaborate on the experimental infrastructure and setup, summarized in Table 9, then we present the methodology used for evaluating NAOMI and comparing both solutions. Further, we identify any limitations to the evaluation and results.
@@ 7.2 Deployment Time
NAOMI supports deployment on a distributed setup, as described in Section 7.1. To test the deployment time of NAOMI on a distributed setup, we automated the deployment using an Ansible playbook. We tested the deployment time from the time the Ansible playbook started executing to the time all the nodes and NAOMI deployments were in a ready state, providing insight into the overhead of a distributed system. To measure the deployment time over multiple runs, we define the average deployment time and its standard deviation as follows:
$T_{\text{dep}}=\bar{T}\pm\sigma_{T},\quad\text{with}$ (1)
$\bar{T}=\frac{1}{N}\sum_{i=1}^{N}T_{i},\quad\sigma_{T}=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}\left(T_{i}-\bar{T}\right)^{2}},$ (2)
where $T_{i}$ is the measured deployment time for the $i$th run, $N$ is the total number of runs, $\bar{T}$ is the average deployment time, and $\sigma_{T}$ is the standard deviation of the average deployment time. We computed the deployment time over $N=10$ measurement runs and plotted the results and the standard deviation utilizing the above defined Eq. 1 and 2.
@@ 7.3 Resource Usage
During the example AI/ML workflow execution, resource utilization was measured at 1 s intervals and CPU usage and RAM usage graphs for both solutions were plotted with timestamps indicating the start of each of the workflow stages. We calculated the average utilization for CPU and RAM during each workflow stage from the measured resource time series. The average utilization for each stage was computed using the following equation across all measurement points for each of the workflow stages:
$\bar{R}_{j}=\frac{1}{M_{j}}\sum_{k=1}^{M_{j}}R_{j,k},$ (3)
where $\bar{R}_{j}$ represents the average resource usage (CPU or RAM) during the $j$th workflow stage, $M_{j}$ is the total number of measurements in stage $j$, and $R_{j,k}$ is the resource usage measured at the $k$th interval.
@@ 7.4 Example AI/ML Workflow Execution Length
We utilized Eq. 1 and Eq. 2 formalized in Section 7.2 to compute the average execution time for each workflow stage and the overall workflow, as well as their corresponding standard deviations. These metrics were calculated across all 5 runs and the results are presented with their respective standard deviation bars in the graphs.
@@ 7.5 Distributed Training
Batch size affects the performance of the distributed training [111], so training time was measured for different batch sizes and compared between different scenarios listed in Section 7.1. For a comparison, we included training time on O-RAN AI/ML Framework with the default scenario. Resource usage metrics and training time were measured in the same manner as described in Section 7.3. We utilized Eq. 1 formalized in Section 7.2 to compute the average training time $\bar{T}_{s}(B)$ as a function of batch size and scenario in Table 9.
@@ 7.6 Inference Latency
As NAOMI supports deployment on a distributed setup, described in Section 7.1, the inference was evaluated for NAOMI when deployed on such setup. We evaluate the ability of load-balancing on multiple nodes and any improvements in mean latency as inference is performed distributively across more CPU cores. On a distributed setup there are 16 cores available for model inference compared to 4 cores on a centralized setup.
To formalize the impact of parallel processing on inference latency, we decompose the overall mean latency $\bar{L}$ into two components: a serial component that includes network communication and other system overhead, and a parallelizable component corresponding to the distributed computation. Referring to Amdahl’s findings [112, 109] this is expressed as:
$\bar{L}=L_{\text{serial}}+\frac{L_{\text{parallel}}}{n},$ (4)
where $L_{\text{serial}}$ represents the portion of work that can not be parallelized such as network communication, scheduling, and other overhead work, $L_{\text{parallel}}$ denotes the latency of tasks that can be executed in parallel, and $n$ is the number of CPU cores available for inference.
The theoretical speedup $S$ when increasing the available cores from 4 to 16 is given by the ratio:
$S=\frac{L_{\text{serial}}+\frac{L_{\text{parallel}}}{4}}{L_{\text{serial}}+\frac{L_{\text{parallel}}}{16}}.$ (5)
In the ideal case where $L_{\text{serial}}$ is negligible compared to $L_{\text{parallel}}$, Eq. 5 simplifies to:
$S\approx\frac{\frac{L_{\text{parallel}}}{4}}{\frac{L_{\text{parallel}}}{16}}=\frac{16}{4}=4,$ (6)
resulting in a theoretical speedup of 4 times.
7.7 Heterogeneous Scale Study
As shown in Section 6 and discussed in Sections 7.5 and  7.6, NAOMI supports distributed training and scalable inference across multiple heterogeneous nodes. To evaluate NAOMI’s scalability under real-world network conditions, we conducted a set of scale study experiments utilizing the devices presented in Section 7.1. Our test environment consisted of an x86_64 virtual machine, three ARM-based Raspberry Pi nodes as presented in Section 7.1, and an additional Intel virtual machine (12 CPU cores, 32 GB RAM) was added to the distributed setup. Nodes were connected through Ethernet cables, creating a representative multi-device distributed and heterogeneous network computing environment.
The scale study increased the number of CPU cores and Ray workers to scale up the computational power of the Ray cluster. We measured QoE model inference performance on parallel requests as discussed in Section 7.6. Model inference is parallelizable and can potentially benefit from a distributed setup as discussed in the preceding sections.
We calculated ideal speedup using Eq. 4, 5, and 6 and measured the actual speedup achieved by increasing the number of CPU cores for model inference. Model replicas were scaled along with CPU cores and Ray workers, each model replica utilizing 1 CPU core. Replicas were distributed along the physical nodes by Ray Serve library and Kubernetes. The measured results and ideal speedup for heterogeneous scale study were plotted on the graph.
@@ 8.2 QoE AI/ML Workflow Execution
Further, the O-RAN solution did not complete the execution of the workflow with the $100x$ dataset, presumably due to timeouts in data extraction stages, which were observed in the logs. NAOMI completed the execution in $1095\pm 20s$ with $100x$ dataset size, which is approximately 3 times longer than the $10x$ dataset. The difference in execution times for NAOMI between the $1x$ and $10x$ datasets is 75 seconds, and between the $10x$ and $100x$ datasets, it is 736 seconds. This indicates that the portion of the workflow execution that takes longer as the dataset size grows increases roughly linearly with the size of the dataset.
@@ 8.3 Distributed Training
From Fig. 11, we can see that NAOMI on a centralized setup executed the QoE workflow in $284\pm 1s$, $359\pm 17s$, and $1095\pm 20s$ for $1x$, $10x$, and $100x$ dataset respectively, while NAOMI on a distributed setup executed the QoE workflow in $356\pm 17s$, $825\pm 37s$, and $5435\pm 539s$ for described dataset sizes. For model training the distributed setup together with distributed model training performs significantly worse. The centralized setup performed $31\%$ faster for $1x$ dataset size, $68\%$ for $10x$ and $83\%$ for $100x$. This shows a significant overhead of distributed model training for this set of training parameters. The distributed setup does not affect data extraction, model deployment, and other (overhead) tasks. In all cases, both setups performed these tasks with a difference of up to 3 seconds, which is within the margin of error.
Distributed model training performing worse than normal training occurs due to the additional communication between nodes and workers in the distributed setup, which increases processing time, particularly with larger datasets and low batch sizes. Synchronizing model updates across the distributed workers introduces delays. As depicted in Fig. 11, when the dataset size increases, we can observe a bigger performance gap between the centralized and distributed setups.
@@ 8.4 QoE Inference Performance
8.5 Heterogeneous Scale Study
Fig. 15 depicts mean inference latency speedup compared to the increasing number of model replicas when running 500 concurrent requests on a deployed QoE model. As discussed in Section 7.7 by increasing the number of replicas a linear latency reduction is expected. The ideal linear speedup is plotted on the graph for comparison.
Figure 15: Inference scale study on distributed heterogeneous nodes.
As we can see from Fig. 15, the speedup increases when increasing the number of replicas. The speedup at 4 replicas is 2.9, and at 8 replicas, it is 8.8. The variations are expected due to heterogeneous architecture, meaning some CPU architectures will perform faster inference on the same amount of CPU cores, so the results will depend on where the models are placed. We can also see that up to 12 replicas, the speedup increases nearly linearly, and at 16, 20, and 24 model replicas, we can observe diminishing returns, as the speedup at 20 replicas is 14.4 and at 24 replicas is 14.9, showing only a slight advantage of having more replicas.
The results show the ability for inference service to scale on NAOMI, however for the chosen model there are no significant benefits of increasing the number of model replicas above 16. Benefits at a larger number of model replicas are expected for models with longer inference latency or for larger amounts of concurrent requests.
@@ 8.6 Workflows Beyond O-RAN
QoE workflow example with $284\pm 1$ second takes, on average, the least time to complete. ImClass and VWW take $670\pm 28s$ and $732\pm 57s$ respectively, while the MNIST workflow example takes $1340\pm 21$ seconds.